Circular No. 9287

Central Bureau for Astronomical Telegrams

Postal Address: Hoffman Lab 209; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbatiau@eps.harvard.edu  ISSN 0081-0304  cbat@iau.org

               URL http://www.cbat.eps.harvard.edu/index.html

Prepared using the Tamkin Foundation Computer Network

THE EDGAR WILSON AWARDS 2014 Under the rules established in 1998 for the Edgar Wilson Awards for

the discovery of comets by amateur astronomers (cf. IAUC 6936), the Awards for 2014 (reckoned with June 11 each year as the annual calendar end date) go to the following individuals: Terry Lovejoy for comet C/2013 R1 (cf. CBET 3649); Michael Schwartz for P/2013 T2 (cf. CBET 3676) and for C/2014 B1 (cf. CBET 3797); Paolo Holvorcem for C/2013 U2 (cf. CBET 3683); Gennadii Borisov for C/2013 V2 (cf. CBET 3694); Vitali Nevski for P/2013 V3 (cf CBET 3695); and Cristovao Jacques for C/2014 E2 (cf. CBET 3828).

In addition, special 2014 Awards outside of the published rules were given by the Smithsonian Astrophysical Observatory to Borisov for comet C/2013 N4 (cf. CBET 3584), which was found using professional instrumentation through his ongoing employment at the Sternberg State Astronomical Institute (his subsequent comet discoveries were made using his own telescopes); to Michel Ory for comet C/2013 V5 (cf. CBET 3713), which was named Oukaimeden; to Jacques, Eduardo Pimental, and Joao Barros for C/2014 A4 (cf. CBET 3783), which was named SONEAR; to Matthias Busch and Rafal Reszelewski for C/2014 C1 (cf. CBET 3800), which was named TOTAS; and to Holvorcem and Schwartz for C/2014 F2 (cf. CBET 3843), which was named Tenagra.

The 2013 Edgar Wilson Awards were announced on IAUC 9269. The Central Bureau learned only in the past week that the Edgar Wilson Awards for 2014 were issued in 2015; there has been no announcement until now.

V5856 SAGITTARII = PNV J18205200–2822100 Further to CBET 4334 and IAUC 9286, A. Pearce (Nedlands, W. Aus-

tralia) reported the following additional magnitudes that showed the nova’s rapid brightening: 2016 Oct. 29.500 UT, 9.9; 31.482, 9.9; Nov. 1.493, 9.1; 2.510, 8.6; 4.486, 8.5; 6.488, 7.6; 7.492, 6.4. S. Nakano, Sumoto, Japan, reported the following CCD magnitudes obtained by K. Yoshimoto with a 180-mm f/2.8 camera lens (and a Nikon D5100 digital camera operating on the G channel): Nov. 2.400, 8.5; 4.389, 8.1; 5.396, 8.1; 7.394, 6.5.

COMET 41P/TUTTLE-GIACOBINI-KRESÁK Visual total-magnitude and coma-diameter estimates reported by J. J.

Gonzalez, Leon, Spain (cf. CBET 4377): 2017 Apr. 1.16 UT, 6.5, 20′ (10×50 binoculars); 7.15, 6.0, 25′ (12×45 monocular); 12.85, 6.2, 17′; 21.11, 6.9, 15′; 28.07, 7.1, 15′.

2017 May 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2017 CBAT  Daniel W. E. Green